Space diary

FALCON FETCH by Kamal Lodaya

[pic: one of Genghis Khan falcon, samurai falcon, surajmal falcon, look under falconry in wikipedia] Falconry has been known in central Asia for at least 3000 years. There are falconers everywhere, in

Japan from the 4th century CE. The pictures [*choose one*] show Mongol ruler Genghis Khan, a Japanese samurai and Maharaja Suraj Mal of Bharatpur in Rajasthan. The peregrine falcon (baaj in Hindi, hayabusa in Chinese and Japanese) is found all over the world, almost as widely as its common prey, pigeons. It was trained by falconers to hunt rabbits, hares, squirrels, small reptiles and birds. Female falcons are twice as large as males, they could hunt larger birds like ducks, swans and fowl, as well as foxes and wolves. Swooping down in dives at 320 km per hour, these are the fastest movers on Earth.

Before the advent of handguns, this was how hunting was done by kings and rich people. Falcons are found in the earliest civilizations, from Mesopotamia, Egypt and China. We know how humans domesticated wolves into dogs. They also domesticated falcons, perhaps from as long ago. Zooarchaeology is a discipline which studies these questions.

In the novels of J.K. Rowling, Harry Potter used Hedwig, an owl, to transport letters. Owls have been used in falconry, although training them is different, since they rely on sound rather than sight. It is carrier pigeons which were used to transport mail, not owls.

Ryugu

[any ryugu pic, preferably showing boulder on pole]

As we reported in Jantar Mantar, the Japanese spacecraft Hayabusa2 reached the 1-kilometre diamondshaped asteroid Ryugu, whose orbit ranges from somewhere close to Earth’s, to somewhere close to the planet Mars’s. Asteroids appear to be all over the solar system. It is now thought there may be some between Venus and Earth also.

Soon after, the spacecraft Osiris Rex (more about it in a later Space diary) reached another diamondshaped asteroid Bennu. Their pictures look alike, though Bennu is smaller. Ryugu’s south pole has a large boulder formation called Otohime Saxum (see the picture). Ryugu and names of its features are from Japanese children’s stories. Ryugu is a palace, and Otohime is a princess.

Hayabusa2’s ”home station” is to remain 20 kilometres from Ryugu. But it can come much closer to within metres of the surface. It first came down and released two Minerva hoppers, later given the names Hibou and Owl, which took pictures. (Hibou is ”owl” in French.) Then it released a lander, called Mascot, which took pictures coming down and sitting on the surface before its batteries ran out.

[www.hayabusa2.jaxa.jp/en/topics/20190328e_{i}llustration/img/T ouchdown1_{a}fter.jpg]

After doing some preliminary trials, Hayabusa2 did a ”touchdown”. A sampling horn hanging down from the spacecraft touched its surface. (The painting is by Akihiro Ikeshita based on images of Ryugu taken by the rovers.) A tantalum bullet was fired into Ryugu, raising a lot of material, some of which fell into the horn. Then the horn was lifted up and the scooped-up sample stowed away in the spacecraft. This will be returned to Earth in 2020, the first large sample of asteroid material. The Hayabusa1 spacecraft earlier only succeeded in bringing back a few dust grains from the asteroid Itokawa.

[www.planetary.org/multimedia/space-images/charts/hayabusa2-conjunction.html]

Meanwhile it so happened that the Earth, moving faster in its orbit than Ryugu, came to a position such that the Sun was between them. This meant that communications between Hayabusa2 and us would stop, so it took a long detour going closer to the Sun and coming back to Ryugu. See the chart. The dates had been calculated using the laws of Johannes Kepler and the equations of Isaac Newton before the spacecraft left Earth.

What is Ryugu made of?

One thing which interests astronomers a great deal is what asteroids like Ryugu and Bennu are made of. Both are carbonaceous chondrites, with rubble-like material similar to meteorites which fall on Earth, with many organic compounds, which have plenty of carbon atoms. The colour spectrum of Ryugu from Earth, from where it is just a point of light like a star, suggested that it may have hydrated compounds, those formed in the presence of water. Perhaps, we thought, Ryugu might have water? Perhaps impacts from asteroids like Ryugu brought water to Earth? Could we bring these compounds back to Earth and check?

The spectrum of Ryugu taken by Hayabusa2 from much closer, also shows the presence of hydrated